Factor by grouping: 3x^3+2x^2+3x+2
Found 2 solutions by user_dude2008, mananth:
Answer by user_dude2008(1862)

3x^3+2x^2+3x+2 = x^2(3x+2)+1(3x+2) = (x^2+1)(3x+2)

Answer by mananth(16949)

3x^3+2x^2+3x+2
..
x^2(3x+2)+1(3x+2)
(3x+2)(x^2+1)
..
m.ananth@hotmail.ca